OpenStudy (anonymous):

Is the point (2, 2 1/2) on the graph of y=x+2? How do you know?

OpenStudy (anonymous):

please help me :(

OpenStudy (solomonzelman):

plg you point into the equation, plug in 2 for x, and 2 1/2 for y.

OpenStudy (solomonzelman):

\(\large\color{black}{ y=x+2 }\) if I had a point (-1,1) I would do this, \(\large\color{black}{ (1)=(-1)+2 }\) \(\large\color{black}{ 1=-1+2 }\) \(\large\color{black}{1=1 }\) So you know that the point (-1,1) is on the line.

OpenStudy (solomonzelman):

you need to do the same exact thing, but with \(\large\color{black}{ (2~,~2\frac{1}{2}) }\)

OpenStudy (anonymous):

oh ok

OpenStudy (solomonzelman):

I would prefer you to do it in here, in this post, so that I can see what you do, but you don't have to...

OpenStudy (anonymous):

ok give me a second

OpenStudy (solomonzelman):

sure

OpenStudy (anonymous):

so 2 1/2=2+2

OpenStudy (solomonzelman):

so far so good..

OpenStudy (anonymous):

the point 2, 2 1/2 is NOT on the graph of y=x+2

OpenStudy (solomonzelman):

Yes, because when you plug it into the equation of the line, the sides are not equal. Correct!

OpenStudy (anonymous):

thank you so much for the help

OpenStudy (solomonzelman):

you really quickly got a hang of it... you welcome !